# CLoQ: Enhancing Fine-Tuning of Quantized LLMs via Calibrated LoRA Initialization

### v1

@@ 3.1.2 Generalized Low-rank Approximation
Llama2-7B Llama2-13B
Method Bit Wiki ($\downarrow$) GSM8K ($\uparrow$) Wiki ($\downarrow$) GSM8K ($\uparrow$)
LoRA 16 5.08 36.9 5.12 45.3
QLoRA 4 5.70 35.1 5.22 39.9
LoftQ 4 5.24 35.0 5.16 45.0
CLoQ 4 5.25 40.6 4.78 49.3
QLoRA 3 5.73 32.1 5.22 40.7
LoftQ 3 5.63 32.9 5.13 44.4
CLoQ 3 5.45 39.9 4.92 48.1
QLoRA 2 N.A. N.A. N.A. N.A.
LoftQ 2 7.85 20.9 7.69 25.4
CLoQ 2 6.51 33.7 5.73 41.7
4 Experiment
@@
Models and Datasets. We test CLoQ on Llama2-7b and Llama2-13b (Touvron et al., 2023) models. Following prior works (Frantar et al., 2022a), we randomly sample 128 instances, each with a context length of 2048 tokens, from the WikiText-2 dataset (Merity et al., 2016) to serve as the calibration set for quantization. Then, we fine-tune and evaluate the models on WikiText-2 for language modeling. For single arithmetic reasoning tasks, we fine-tune and evaluate on the GSM8K (Cobbe et al., 2021). For multi arithmetic reasoning, we fine-tune the models on Math10K (Hu et al., 2023) and then evaluate the test sets of AQuA (Ling et al., 2017), GSM8K, MAWPS (Koncel-Kedziorski et al., 2016) and SVAMP (Patel et al., 2021). For commonsense reasoning tasks, we fine-tune the models on Commonsense170K (Hu et al., 2023) and evaluate on eight representative tasks: BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018) and OBQA (Mihaylov et al., 2018).
Baselines. We compare with LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2023), GPTQ-LoRA (GPT, 2023) and LoftQ (Li et al., 2023). LoRA is often considered as the benchmark for fine-tuning performance. QLoRA incorporates NF-quantization, with its low-rank initialization aligning with the standard LoRA method. GPTQ-LoRA integrates OPTQ for the base weights and fine-tunes the LoRA component while preserving its low-rank initialization as in the standard LoRA, with the quantized weights kept frozen. In contrast, LoftQ and LQ-LoRA carefully initialize the quantized weight and low-rank matrices by solving some optimization problems to minimize the approximation error.
4.1 Implementation details
Quantization. We quantize the weights of all linear layers in the base model using MagR preprocessing (Zhang et al., 2024a) followed by OPTQ (Frantar et al., 2022a). The quantization scheme employs uniform (a.k.a. INT) and asymmetric quantization, with a default group size of 64. After quantization, we compute the LoRA components $\boldsymbol{A}$ and $\boldsymbol{B}$, which are maintained in FP16 precision.
Fine-tuning. Following prior works (Dettmers et al., 2023; Li et al., 2023; Liao et al., 2024), we fine-tune the models using the standard LoRA configuration, with modifications to the LoRA initialization and learning rates. The quantized weights remain fixed, and only the LoRA adapter matrices are trainable during fine-tuning. The rank of the LoRA adapters is consistently set to 64 across all methods. For optimization, we use AdamW (Loshchilov, 2017). All experiments are conducted on NVIDIA A100 GPUs with 80GB of memory.
4.2 Fine-tuning Results
Language modeling. We evaluate the models by reporting the perplexity of language generation on WikiText-2. As shown in Table 2, CLoQ consistently delivers the best performance across the majority of bit levels when compared to other methods. Notably, at INT2, CLoQ demonstrates its effectiveness by achieving a perplexity improvement of 1.34 on Llama2-7B over LoftQ. This result highlights CLoQ’s ability to maintain superior performance even under ultra-low bit quantization constraints.
Arithmetic reasoning (single task). To evaluate the models on GSM8K, we extract numerical answers from the generated solutions and determine accuracy by analyzing these extracted values. As shown in Table 2, CLoQ achieves better performance across different model sizes and quantization bit levels. CLoQ reaches an accuracy of $33.7\%$ at INT2 on Llama2-7B, which surpasses that of LoRA at INT4 across all model sizes. More remarkably, the accuracy of INT2 CLoQ on the Llama2-13B model outperforms that of INT4 QLoRA. On the Llama2-7B model, the accuracy of INT2 CLoQ is better than that of INT3 QLoRA and is comparable with INT4 QLoRA.
Arithmetic reasoning. To evaluate CLoQ’s effectiveness across multiple arithmetic reasoning tasks, we fine-tuned models on the Math10K dataset and assessed their performance on four separate math reasoning benchmarks, demonstrating their adaptability to diverse mathematical challenges. As shown in Table 1, CLoQ consistently outperforms other methods across various model sizes and quantization levels, achieving superior accuracy both on average and on individual datasets. Notably, at INT2 quantization, CLoQ delivers substantial gains, improving accuracy by 3.6% and 4.8% over LoftQ on Llama2-7B and Llama2-13B, respectively. Even at INT4, CLoQ surpasses both LoRA and QLoRA on Llama2-7B. Moreover, for complex reasoning tasks such as GSM8K and SVAMP, CLoQ achieves 5.2% and 6.2% higher accuracy than LoftQ on Llama2-7B, highlighting its robustness in challenging problem settings. These results underscore CLoQ’s remarkable potential in handling intricate reasoning tasks with enhanced accuracy, even under ultra-low-bit quantization.
Table 4: Finetuning results of WikiText and GSM8K for g128 and per-channel quantization.
Llama2-7B Llama2-13B
Method Bit Wiki ($\downarrow$) GSM8K ($\uparrow$) Wiki ($\downarrow$) GSM8K ($\uparrow$)
LoRA 16 5.08 36.9 5.12 45.3
LoftQ 4 5.24 35.0 5.16 45.0
CLoQ 4 5.28 39.5 4.79 49.3
LoftQ 3 5.63 32.9 5.13 44.4
CLoQ 3 5.55 38.7 4.99 46.9
3g128 5.48 39.0 4.94 48.6
LoftQ 2 7.85 20.9 7.69 25.4
CLoQ 2 6.95 31.6 6.07 40.0
2g128 6.77 32.1 5.90 40.4
Table 5: Accuracy on four arithmetic reasoning tasks for g128 and per-channel quantization. The LoRA rank is $r=64$.
Llama2-7B Llama2-13B
LoRA 16 43.6 59.4 85.0 27.0 53.7 55.3 67.7 87.4 24.4 58.7
CLoQ 4 42.8 59.5 88.2 25.2 53.9 53.7 67.2 89.1 26.8 59.2
CLoQ 3 41.4 58.5 86.6 24.4 52.7 51.4 65.2 88.2 27.6 58.1
3g128 42.2 58.9 89.1 24.0 53.6 52.8 64.7 89.5 27.2 58.3
CLoQ 2 34.0 49.1 87.0 24.8 48.7 40.5 55.2 88.7 28.0 53.1
2g128 33.9 46.9 85.7 26.4 48.9 41.4 58.0 89.1 25.2 53.4
At INT2, CLoQ delivers a substantial performance boost over LoftQ, with an average accuracy improvement exceeding 10%, even approaching the performance levels typically observed with INT4 QLoRA. For instance, INT2 CLoQ exhibits only a minimal average accuracy drop of 0.04% compared to INT4 QLoRA. Additionally, at INT4, CLoQ attains impressive accuracy, reducing the gap to FP16 LoRA to just 0.4%. These results indicate that CLoQ significantly enhances LoRA’s learning capacity, enabling it to better adapt to a diverse range of tasks.
@@ 4.3 Ablation study
Quantization with different group sizes. We examine the performance of CLoQ for group-wise-128 (g128) and per-channel quantization. Table 4 shows the results for Wikitext-2 and GSM8K. Table 5 shows the result on multiple arithmetic reasoning tasks. CLoQ demonstrates superior performance even under per-channel quantization, achieving results remarkably close to those of g64 quantization. Moreover, it consistently outperforms LoftQ in these settings, demonstrating its effectiveness across all quantization bit levels.
@@
$(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}}\boldsymbol{H}_{\text{had}},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}}\boldsymbol{H}_{\text{had}})$ 2 6.51 29.0
@@
LoRA initialization with different $(\boldsymbol{A},\boldsymbol{B})$ combinations. Furthermore, we investigate the performance of various combinations of ($\boldsymbol{A},\boldsymbol{B}$) in Algorithm 1, as shown in Table 6. We tried four different combinations of $(\boldsymbol{A},\boldsymbol{B})$, including $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r})$, $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}})$, $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}}\boldsymbol{H}_{\text{had}},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}}\boldsymbol{H}_{\text{had}})$ with $\boldsymbol{H}_{\text{had}}$ being the order-$r$ ($r=64$) Hadamard matrix, and the default choice $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r},\,\boldsymbol{V}_{:r})$. Table 6 shows that the default combination of initialized adapters gives the best performance in the subsequent fine-tuning phase.
@@ 5 Related Work
When quantizing pre-trained models, QLoRA (Dettmers et al., 2023) primarily emphasizes the quantization process, often overlooking the critical importance of subsequent LoRA fine-tuning. It adopts the fixup initialization strategy used in LoRA, attaching zero-initialized low-rank adapters to the quantized pre-trained model. However, the discrepancies introduced by quantization, especially in extremely low-bit regimes, can significantly impact the initialization of LoRA fine-tuning, ultimately affecting the overall fine-tuning performance. LoftQ (Li et al., 2023) jointly optimizes the quantized weights $\boldsymbol{Q}$ and the low-rank adapter matrices $\boldsymbol{A}$ and $\boldsymbol{B}$ by solving the following optimization problem:
@@
LQ-LoRA (Guo et al., 2024b) assigns importance weights to each parameter by evaluating its sensitivity to output variations, thereby guiding the decomposition process toward critical regions. Specifically, the Fisher matrix is used to weight the reconstruction objective during decomposition. To solve the resulting weighted SVD problem, LQ-LoRA assumes row and column homogeneity in the Fisher matrix, allowing the use of standard SVD techniques at the cost of theoretical precision. However, this approach has limitations. It relies on approximations rather than solving the weighted SVD problem exactly, which may lead to suboptimal results. Furthermore, computing the Fisher matrix requires back-propagation through the pre-trained model, introducing additional computational overhead. Similar to our work, ApiQ (Liao et al., 2024) also employs an activation-aware initialization strategy utilizing calibration data. However, it relies on two activation matrices—one obtained from the pre-trained model and the other from the quantized model, whereas CLoQ uses only a single pre-trained activation matrix. ApiQ optimizes the discrepancy using standard back-propagation, whereas CLoQ adopts a fully gradient-free approach. By avoiding the computational overhead of back-propagation, CLoQ enables faster adaptation while maintaining high performance across various tasks.
@@ 6 Concluding Remarks
Acknowledgements
This work was partially supported by NSF grants DMS-2208126, DMS-2110836, IIS-2110546, SUNY-IBM AI Research Alliance Grant, and a start-up grant from SUNY Albany. We would also like to thank SUNY Albany for providing access to the Nvidia A100 GPUs.
@@ A.2 Arithmetic reasoning
Following the framework proposed by (Hu et al., 2023), we adopt a more integrated approach by training a single model across multiple tasks. Specifically, we fine-tune Llama2-7B and Llama2-13B on Math10K, a dataset that aggregates training samples from GSM8K, MAWPS, MAWPS-single, and AQuA. After finetuning, the models are tested on the evaluation sets of AQuA, GSM8K, MAWPS, and SVAMP. The hyper-parameters used for fine-tuning are detailed in Table 7 and Table 8. Moreover, instead of conducting evaluations at every epoch, we assess model performance only after the final epoch.
@@ A.3 Commonsense reasoning
To evaluate the commonsense reasoning capabilities of CLoQ, we consider eight key benchmark tasks: BoolQ, PIQA, SIQA, HellaSwag, WinoGrande, ARC-e, ARC-c, and OBQA. We adopt the framework proposed by (Hu et al., 2023) and fine-tune a single model across all these tasks instead of training separate models. We fine-tune Llama2-7B and Llama2-13B on the merged training set and measure accuracy on the corresponding test sets. The hyper-parameters used for fine-tuning are detailed in Table 7 and Table 8. For evaluation, we forgo per-epoch assessments and instead report the final model’s performance after the last epoch.

### v2

@@ 4.2 Fine-tuning Results
Table 4: Accuracy on four arithmetic reasoning tasks. The LoRA rank $r$ is 64 for all methods. CLoQ accuracies are averaged over 5 random runs, with standard deviations reported.
Llama3-8B
Method Bit GSM8K SVAMP MAWPS AQuA Avg. $\uparrow$
LoRA 16 68.8 76.1 90.3 31.5 66.7
LoftQ 2 35.6 52.1 87.0 25.2 50.0
ApiQ-bw 2 47.0 67.2 88.2 27.2 57.4
CLoQ 2 $50.7_{\pm 0.35}$ $67.5_{\pm 0.86}$ $88.3_{\pm 0.91}$ $29.5_{\pm 0.56}$ $59.0_{\pm 0.34}$
At INT2, CLoQ delivers a performance boost over ApiQ-bw, with an average accuracy improvement exceeding 0.7%, even approaching the performance levels typically observed with INT4 QLoRA. For instance, INT2 CLoQ exhibits only a minimal average accuracy drop of 0.04% compared to INT4 QLoRA. Additionally, at INT4, CLoQ attains impressive accuracy, reducing the gap to FP16 LoRA to just 0.4%. These results indicate that CLoQ significantly enhances LoRA’s learning capacity, enabling it to better adapt to a diverse range of tasks.
@@
Fine-tuning on mixed dataset. We evaluate the performance of CLoQ for LoRA fine-tuning on a mixed dataset comprising Math10K and 5K samples from a Commonsense dataset, and assess accuracy on four arithmetic reasoning tasks. Interestingly, we find that fine-tuning on the mixed dataset leads to a drop in arithmetic reasoning accuracy compared to fine-tuning solely on the Math10K dataset, as shown by Table 6. For example, for 4-bit Llama2-7B, the averaged accuracy of CLoQ fine-tuned on mixed dataset drops to 51.2% from 54.6% (in Table 3). However, CLoQ still consistently outperforms LoftQ in this setting across different bit-widths.
Table 6: Accuracy of Llama2-7B for four arithmetic reasoning finetuned on mixed dataset.
Arithmetic reasoning
Method Bit GSM8K SVAMP MAWPS AQuA Avg. $\uparrow$
LoftQ 4 36.8 55.3 83.2 24.8 50.0
CLoQ 4 38.8 54.9 85.7 25.2 51.2
LoftQ 2 21.8 40.7 74.4 24.0 40.2
CLoQ 2 27.2 47.0 80.3 24.4 44.7
@@ 4.3 Ablation study
LoRA initialization with different $(\boldsymbol{A},\boldsymbol{B})$ combinations. Furthermore, we investigate the performance of various combinations of ($\boldsymbol{A},\boldsymbol{B}$) in Algorithm 1, as shown in Table 7. We tried three different combinations of $(\boldsymbol{A},\boldsymbol{B})$, including $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r})$, $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}},\,\boldsymbol{V}_{:r}\boldsymbol{\Sigma}_{:r}^{\frac{1}{2}})$, and the default choice $(\boldsymbol{R}^{-1}\boldsymbol{U}_{:r}\boldsymbol{\Sigma}_{:r},\,\boldsymbol{V}_{:r})$. Table 7 shows that the default combination of initialized adapters gives the best performance in the subsequent fine-tuning phase.
@@
Calibration data size. We explore the sensitivity of CLoQ to the size of the calibration dataset. As shown in Table 8, CLoQ exhibits strong robustness to the calibration size. Across both INT4 and INT2, the overall performance remains consistently stable as the calibration size varies from 32 to 256. A calibration dataset of 128 samples, commonly used as the default in PTQ, yields slightly better results. These findings indicate that CLoQ does not heavily depend on the specific choice of calibration set size, which enhances its practicality and ease of deployment in real-world scenarios.
Table 8: Accuracy of different calibration dataset sizes for Llama2-7B.
Arithmetic reasoning
Calibration Size Bit Wiki. $\downarrow$ GSM8K $\uparrow$ GSM8K SVAMP MAWPS AQuA Avg. $\uparrow$
32 4 5.34 40.2 43.4 58.7 87.4 28.0 54.4
64 4 5.33 40.9 44.7 57.6 87.0 25.2 53.6
128 (default) 4 5.25 40.6 43.6 60.3 87.0 27.6 54.6
256 4 5.33 40.1 42.6 59.6 87.0 26.0 53.8
32 2 6.62 32.5 35.5 50.5 86.6 24.0 49.1
64 2 6.56 33.5 35.5 54.0 84.5 25.6 49.9
128 2 6.51 33.7 34.7 52.0 86.1 24.1 49.2
256 2 6.49 33.2 33.5 47.2 87.4 27.6 48.9
Sequence length. Additionally, we present the results for different sequence lengths when finetuning the 2-bit Llama2-7B model in Table 9 on arithmetic reasoning tasks. It shows that the fine-tuning accuracy nicely improves as the sequence length increases.
Table 9: Accuracy of finetuned 2-bit Llama2-7B for four arithmetic reasoning with different sequence lengths.
Arithmetic reasoning
Sequence length GSM8K SVAMP MAWPS AQuA Avg. $\uparrow$
256 35.0 50.7 87.0 22.1 48.7
512 34.7 52.0 86.1 24.1 49.2
1024 34.1 52.8 87.8 24.0 49.7
2048 34.5 52.9 87.8 24.6 50.0
Initialization Cost and Latency/Memory Benefits. We compare the duration and GPU memory usage of Initialization between CLoQ and other baseline methods. As shown in Table 10, CLoQ achieves efficient and scalable initialization, offering notable reductions in both runtime and memory consumption. For instance, CLoQ requires only 0.7 hours and 9GB of memory on LLaMA2-7B, outperforming ApiQ-lw (4.1h) and ApiQ-bw (1.3h) by large margins. On the larger LLaMA2-13B, CLoQ remains competitive, requiring just 1.5 hours and 13GB of memory, compared to 6.5 hours (ApiQ-lw) and 27GB (LoftQ). These results demonstrates CLoQ’s strong practicality for large-scale model under limited hardware budgets.
Table 10: The duration and peak GPU memory used for Llama2.
Size Method Duration Peak GPU memory
Llama2-7B LoftQ 0.6h 14GB
ApiQ-lw 4.1h 6GB
ApiQ-bw 1.3h 12GB
CLoQ 0.7h 9GB
Llama2-13B LoftQ 1.3h 27GB
ApiQ-lw 6.5h 9GB
ApiQ-bw 2.4h 17GB
CLoQ 1.5h 13GB
@@ 5 Related Work
When quantizing pre-trained models, QLoRA Dettmers et al. (2023) primarily emphasizes the quantization process, often overlooking the critical importance of subsequent LoRA fine-tuning. It adopts the fixup initialization strategy used in LoRA, attaching zero-initialized low-rank adapters to the quantized pre-trained model. However, the discrepancies introduced by quantization, especially in extremely low-bit regimes, can significantly impact the initialization of LoRA fine-tuning, ultimately affecting the overall fine-tuning performance. LoftQ Li et al. (2023a) jointly optimizes the quantized weights $\boldsymbol{Q}$ and the low-rank adapter matrices $\boldsymbol{A}$ and $\boldsymbol{B}$ by solving the following optimization problem:
@@
LQ-LoRA Guo et al. (2024b) assigns importance weights to each parameter by evaluating its sensitivity to output variations, thereby guiding the decomposition process toward critical regions. Specifically, the (diagonal) Fisher matrix is used to weight the reconstruction objective during decomposition. To solve the resulting weighted SVD problem, LQ-LoRA assumes row and column homogeneity in the Fisher matrix, allowing the use of standard SVD techniques at the cost of theoretical precision. However, this approach has limitations. It relies on approximations rather than solving the weighted SVD problem exactly, which may lead to suboptimal results. Furthermore, computing the Fisher matrix requires back-propagation through the pre-trained model, introducing additional computational overhead. Similar to our work, ApiQ Liao et al. (2024) also employs an activation-aware initialization strategy utilizing calibration data. However, it relies on two activation matrices—one obtained from the pre-trained model and the other from the quantized model, whereas CLoQ uses only a single pre-trained activation matrix. ApiQ optimizes the discrepancy using standard back-propagation, whereas CLoQ adopts a fully gradient-free approach. By avoiding the computational overhead of back-propagation, CLoQ enables faster adaptation while maintaining high performance across various tasks.
@@ 6 Concluding Remarks
7 Acknowledgments
This work was partially supported by NSF grants DMS2208126, DMS-2110836, IIS-2110546, SUNY-IBM AI Research Alliance Grant, UAlbany-IBM CEAIS grant, and a start-up grant from SUNY Albany. We would also like to thank SUNY Albany for providing access to the Nvidia DGX Cloud.
Appendix
@@ A.2 Arithmetic reasoning
Following the framework proposed by Hu et al. (2023), we adopt a more integrated approach by training a single model across multiple tasks. Specifically, we fine-tune Llama2-7B and Llama2-13B on Math10K, a dataset that aggregates training samples from GSM8K, MAWPS, MAWPS-single, and AQuA. After finetuning, the models are tested on the evaluation sets of AQuA, GSM8K, MAWPS, and SVAMP. The hyper-parameters used for fine-tuning are detailed in Table 11 and Table 12. Moreover, instead of conducting evaluations at every epoch, we assess model performance only after the final epoch.
@@ A.3 Commonsense reasoning
To evaluate the commonsense reasoning capabilities of CLoQ, we consider eight key benchmark tasks: BoolQ, PIQA, SIQA, HellaSwag, WinoGrande, ARC-e, ARC-c, and OBQA. We adopt the framework proposed by Hu et al. (2023) and fine-tune a single model across all these tasks instead of training separate models. We fine-tune Llama2-7B and Llama2-13B on the merged training set and measure accuracy on the corresponding test sets. The hyper-parameters used for fine-tuning are detailed in Table 11 and Table 12. For evaluation, we forgo per-epoch assessments and instead report the final model’s performance after the last epoch.